Keller Cosmetics maintains an operating profit margin of 4% and asset turnover ratio of 5 what is its ROA? If its debt equity ratio is 1, interest payments and taxes are $9,800, and EBIT is $29,000, what is its ROE?
Answer #1

ROA = Operating Profit Margin * Asset Turnover ratio =4%*5=20%

Debt Equity Ratio is 1 so Total assets to equity ratio =2:1
Net Income =EBIT - Interest - Taxes =29000-9800-9800 =9400
Operating income after taxes =29000-9800 =192000

ROE=Assets to equity ratio* Net Income/Operating income after taxes * ROA =2*(9400/192000)*20% =19.58%